OpenStudy (anonymous):

csc x cot x (1-cos^2 x) = ______ x @terenzreignz :)

OpenStudy (anonymous):

idk this one :( no answer choices! :/

terenzreignz (terenzreignz):

Pythagorean identity is key here, along with a few more tricks... \[\Large \cos^2(x) + \sin^2(x) = 1\]

OpenStudy (anonymous):

okay :) how do i use that to solve? :/

terenzreignz (terenzreignz):

Magic.

OpenStudy (anonymous):

lol :P i love magic! :P

terenzreignz (terenzreignz):

If we bring \(\large \cos^2(x)\) to the other side, we get \[\Large \color{blue}{\sin^2(x) }= \color{red}{1-\cos^2(x)}\]

OpenStudy (anonymous):

okay :) i see what you did there!

terenzreignz (terenzreignz):

In that respect, this \[\Large \csc(x)\cot(x)(\color{red}{1-\cos^2(x)})\] becomes...?

OpenStudy (anonymous):

csc(x)cot(x)(sin^2(x)) ?

terenzreignz (terenzreignz):

yes, that's right. Now, let's break down the fancy csc and cot functions into simpler forms, and also, let's separate sin^2(x) into sin(x)sin(x) \[\Large \frac{1}{\sin(x)}\cdot\frac{\cos(x)}{\sin(x)}\cdot\sin(x)\cdot\sin(x)\]

OpenStudy (anonymous):

okay :)

terenzreignz (terenzreignz):

Now... I'll just take my sword and do some slashing... \[\Large \frac{1}{\cancel{\sin(x)}}\cdot\frac{\cos(x)}{{\sin(x)}}\cdot\cancel{\sin(x)}\cdot{\sin(x)}\]

terenzreignz (terenzreignz):

And again...\[\Large \frac{1}{\cancel{\sin(x)}}\cdot\frac{\cos(x)}{\cancel{\sin(x)}}\cdot\cancel{\sin(x)}\cdot\cancel{\sin(x)}\]

OpenStudy (anonymous):

hahaha thee sword is thy weapon :P

terenzreignz (terenzreignz):

I use a Chinese Jian. And now, we are left with \[\huge \cos(x)\]

OpenStudy (anonymous):

lol no clue what that is, but ima guess its some kind of super cool sword lol :P okay! so the answer is just "cos" right? "cos" goes into the blank yeah? :)

OpenStudy (anonymous):

okay haha googled it.. so its a sword that looks like it needs to be sheathed hahah :P cooool :)

terenzreignz (terenzreignz):

Well, you have your answer :P

OpenStudy (anonymous):

awesome!! :) yay! thankS!!! :D